# Transfer Learning for T-Cell Response Prediction

### v1

@@ Abstract
Motivation: We study the prediction of T-cell response for specific given peptides, which could, among other applications, be a crucial step towards the development of personalized cancer vaccines. It is a challenging task due to limited, heterogeneous training data featuring a multi-domain structure; such data entail the danger of shortcut learning, where models learn general characteristics of peptide sources, such as the source organism, rather than specific peptide characteristics associated with T-cell response.
Results: Using a transformer model for T-cell response prediction, we show that the danger of inflated predictive performance is not merely theoretical but occurs in practice. Consequently, we propose a domain-aware evaluation scheme. We then study different transfer learning techniques to deal with the multi-domain structure and shortcut learning. We demonstrate a per-source fine tuning approach to be effective across a wide range of peptide sources and further show that our final model outperforms existing state-of-the-art approaches for predicting T-cell responses for human peptides.
Availability: Source code for all studies, compiled data, and final model are available under github.com/JosuaStadelmaier/T-cell-response-prediction
Contact: josua.stadelmaier@uni-tuebingen.de
@@ 1 Introduction
The human immune system consists of humoral and cell-mediated mechanisms, with T cells playing a significant role in the latter. T cells recognize and eliminate cancerous or infected cells by detecting peptides presented on the cell surface by major histocompatibility complex (MHC) molecules, which can be grouped into two classes, MHC I and MHC II (Murphy and Weaver, 2017).
@@
Peptide-based vaccines are a promising approach for the personalized treatment of cancer (Sahin and Türeci, 2018) and might allow more precise vaccines against infectious diseases (Heitmann et al., 2022). Such vaccines can be developed by selecting peptides that are specific to a pathogen or tumor. Since only a limited number of peptides can be included in the vaccine, their selection should be based on their probability of inducing a T-cell response. Predicting this probability is an important task in the development of peptide vaccines (Sahin and Türeci, 2018).
@@
From a computational perspective, this task can be separated into two sub-tasks: first predicting whether a peptide is presented on the cell surface by an MHC molecule and afterwards predicting whether a presented peptide induces a T-cell response. For the first task, a large amount of available experimental data has led to accurate machine learning-based predictions (Reynisson et al., 2020; O’Donnell et al., 2020; Cheng et al., 2021). The second task can, despite some attempts (Gfeller et al., 2023; Dhanda et al., 2018), still be considered as an open problem, in part due to fewer experimental data from T-cell assays (Vita et al., 2018).
@@
Aside from small sample sizes, another challenge arises from peptides in T-cell response data originating from different sources, such as viruses, bacteria, or human proteins. Moreover, peptides feature patterns that are specific to the allele of the MHC molecule they are presented by (Rammensee et al., 1995). Both factors lead to data heterogeneity, which is either unaccounted for by existing approaches (Dhanda et al., 2018; Ogishi and Yotsuyanagi, 2019) or handled by limiting the model to selected MHC alleles and considering small peptides, such that positions that are affected by MHC allele-specific patterns can be identified (Gfeller et al., 2023; Calis et al., 2013).
@@
To leverage all the available data, we investigate more flexible approaches that do not limit the selection of peptides based on their source, length, and presenting MHC allele. Based on our data analysis, we show how the heterogeneity of the resulting data set can be viewed as a multi-domain structure, which motivates the application of different transfer learning techniques in combination with the flexible transformer architecture (Vaswani et al., 2017). We further investigate how the prediction performance on individual domains is affected by including data from the other domains in the training process, where improved performance is denoted as positive transfer and reduced performance as negative transfer (Rosenstein, 2005).
@@
For all empirical studies it is critical to keep in mind that having several underlying domains with varying fractions of positives entails the risk of shortcut learning (Geirhos et al., 2020), that is, using domain-specific features to classify peptides based on their domain identity instead of T-cell response-specific patterns. To obtain shortcut-invariant model performance estimates, we thus propose a domain-aware evaluation scheme.
@@
Our results indicate that shortcuts based on peptide sources and MHC alleles are indeed learned by the transformer and lead to inflated performance estimates when the domain-aware evaluation is not applied. We further observe that per-source fine-tuning is effective in enabling positive transfer across a wide range of peptide sources and show that our final model outperforms existing state-of-the-art approaches for predicting T-cell responses for human peptides.
@@ 2.1 Construction of a T-cell response data set
For some peptides in the data set, the MHC alleles are not specified in the IEDB. Additionally, the MHC allele information is incomplete since one peptide can often be presented on several MHC alleles. However, not all combinations of peptides and MHC alleles are tested. Since the existing MHC information in the IEDB is mostly based on predictions, we use existing models to obtain a list of MHC alleles for each peptide in a consistent way. NetMHCpan 4.1 (Reynisson et al., 2020) serves as a predictor for peptides with MHC class I alleles and NetMHCIIpan 4.0 (Reynisson et al., 2020) for peptides with MHC class II alleles. We select the 100 most frequent alleles, which account for 98.7% of peptide-MHC combinations in the IEDB data set, and use all MHC class-matched combinations of peptides and MHC alleles as input for the two predictors.
@@ 2.2 Domain structure in T-cell response data
Motivated by the size of the binding core in peptides (Nielsen et al., 2007), we define that peptides correspond to the same cluster if they share a sub-sequence of length nine. Using this definition, the Sankey diagram in Figure 1c confirms that the seven largest peptide clusters typically consist of peptides from only a few sources.
@@
The distinct binding motifs of MHC alleles (Rammensee et al., 1995) introduce an MHC-based domain structure in T-cell response data. We visualize the sequence logos for two exemplary MHC alleles and peptides of length nine, separately for T-cell response positives and negatives, in Figure 1d. At the anchor positions two and nine, there are obvious differences between the MHC alleles in the sequence logos. However, comparing sequence logos between T-cell response positives and negatives does not reveal obvious differences.
@@ 2.3 Transformer model
We use the transformer architecture (Vaswani et al., 2017), as shown in Figure 2, to capture T-cell response-specific patterns in peptide sequences. To apply the transformer, we formalize the data set to consist of data points $\{(\bm{x}_{i},y_{i})\}_{i=1}^{N_{\text{data}}}$, where $\bm{x}_{i}=(x_{i,1},\dots,x_{i,L})$ is the $i$-th peptide sequence with each of the $x_{i,l}$ representing one of 20 amino acids or a padding token. The padding token allows the sequences to all have the same length $L$, which corresponds to the maximal peptide length. We denote the binary label indicating whether the peptide $\bm{x}_{i}$ leads to a T-cell response with $y_{i}$.
Our model gets a one-hot representation of the peptide sequence $\bm{x}_{i}$ as input. Then the model maps each amino acid or padding token $x_{i,j}$ of the peptide sequence to a learned embedding $\bm{e}_{i,l}\in\mathbb{R}^{d/2}$. To tell the transformer the positions at which the amino acids occur in the peptide, we define a learned embedding $\bm{p}_{l}\in\mathbb{R}^{d/2}$ for each position $l\in\{1,\dots,L\}$. Additionally, we use positional encodings $\bm{s}_{l}\in\mathbb{R}^{d/2}$ based on sine functions with varying frequencies (see Vaswani et al. (2017) for more details). We then add the two variants of positional vectors and concatenate them with the amino acid embeddings to obtain a vector representation $\left(\begin{smallmatrix}\bm{e}_{i,l}\\ \bm{p}_{l}+\bm{s}_{l}\end{smallmatrix}\right)\in\mathbb{R}^{d}$ for each element $x_{i,l}$ of the peptide sequence $\bm{x}_{i}$.
The resulting sequence of vector representations is the input of a transformer encoder layer, which updates the vector representations based on interactions between amino acids using a mechanism called multi-head self-attention (see Vaswani et al. (2017) for more details).
As the output of the transformer encoder is again a sequence of vector representations, we apply mean pooling to obtain a single vector $\bm{h}_{i}\in\mathbb{R}^{d}$, which is a latent representation of the $i$-th peptide. In a similar way as described here, the transformer architecture has been used in earlier work to learn representations of peptides (Cheng et al., 2021) and proteins (Rao et al., 2019).
@@ 2.4 Adversarial domain adaptation
To address these issues, we apply a transfer learning technique called adversarial domain adaptation (Tzeng et al., 2017), to encourage the model to perform predictions independently of the domain identity of a peptide. Unlike the typical domain adaptation setting with a large amount of training data for a source domain and no or little training data for the target domain (Pan and Yang, 2009), in our setting, there are more than two domains and all domains can act as target and source domain at the same time. We use a variant of adversarial domain adaptation that is adapted to our setting. This variant is very similar to multi-domain adversarial learning (Schoenauer Sebag et al., 2019).
@@ 2.5 Per-source fine-tuning
When one machine learning model is trained on data from several domains, the performance on individual domains can be reduced compared to models that are trained on data from one domain only. Such instances of negative transfer are more likely to occur if domains are too dissimilar (Rosenstein, 2005). Negative transfer is also possible when applying adversarial domain adaptation because there is only one model for all domains. Additionally, data points from all domains are encouraged to be represented similarly in the model. To avoid negative transfer, a transfer learning technique is needed that provides more flexibility to adapt to individual domains, while still allowing positive transfer from other domains.
As an alternative to adversarial domain adaptation, we propose to pre-train a T-cell response model until convergence on all domains and then continue training on only one domain. This second step is also called fine-tuning. We allow all model parameters to be updated during fine-tuning such that the model has the necessary flexibility to correct instances of negative transfer. This transfer learning technique is commonly applied in combination with the transformer and proteins (Rao et al., 2019) or peptides (Cheng et al., 2021).
@@ 2.6 Bag of amino acids baseline
To allow a direct comparison of our proposed transformer-based models to a simpler model that is trained on the same data set, we define the bag of amino acids (Bag-Of-AA) baseline, which can be seen as a simplified version of PRIME (Gfeller et al., 2023).
The Bag-Of-AA baseline counts how often each of the 20 amino acids occurs in a given peptide and puts the frequencies into a 20-dimensional vector. An MLP with one hidden layer receives these vectors as input and is trained to predict the T-cell responses. As with the transformer models, we use the binary cross-entropy as training objective. The Bag-Of-AA model does not consider positional information, that is, it does not see at which positions the amino acids occur in the peptide. The amino acid frequencies allow the model to capture possible preferences for certain amino acids that T cells might have.
@@ 2.7 Model evaluation
Within a peptide cluster, sequences can be nearly identical (cf. Section 2.2). Splitting the data completely at random could lead to such near identical sequences being distributed among training, validation, and test sets. This can potentially cause inflated generalization estimates. To prevent this, we assign peptide clusters as a whole (instead of single peptides) to a randomly selected partition (Nielsen et al., 2007).
@@ 2.8 Detection of shortcut learning
In our analysis, we evaluate model performance to account for such shortcuts. In particular, we group the evaluation data into subsets such that each subset contains peptides from only one combination of peptide source and MHC allele. We further weight by subset size and peptide:MHC assignments to ensure each peptide contributes equally to the overall performance (see Supplement S2 for details).
Since this evaluation setting ensures that performance estimates are not inflated by the use of shortcuts, we refer to this strategy as shortcut adjusted evaluation.
@@ 3 Results
Figure 3: Shortcut learning and the effect of adversarial domain adaptation. The left column shows results for BASE-T and the right column shows corresponding results for ADA-T with adversarial domain adaptation being applied on peptide sources. (a) Model performance on validation data with different settings of accounting for shortcuts in the evaluation. For the "allele adjusted" performance, peptides are grouped by MHC alleles in the evaluation to detect shortcuts based on MHC alleles. "Source adjusted" is analogous for grouping peptides by their source. (b) Distribution of prediction scores, separated by the T-cell response labels. For each label, two distributions are shown that correspond to the two most frequent peptide sources, Human betaherpesvirus 6B (majority of labels is positive) and Vaccinia virus (majority of labels is negative). The markers on the x-axis show the mean prediction scores of the two distributions. (c) t-SNE visualization of the latent peptide representations $\mathbf{h}$.
@@ 3.1 Shortcut learning in practice
We investigate the prevalence of shortcut learning in T-cell response prediction models trained on the full T-cell response data set, that is, on all peptide sources and MHC alleles (MHC I+II) combined. For this purpose, we use the shortcut detection method described in Section 2.8. In order to study whether shortcut learning can be prevented by encouraging domain invariance in the model architecture and whether this constraint could serve as regularizer to improve the model performance, we compare two different models: BASE-T, the baseline transformer of Section 2.3, and ADA-T, the adversarial domain adaptation approach of Section 2.4 with domain adaptation strength set to $\lambda=10$. The results for both models are summarized side by side in Figure 3.
@@
The prediction performance in terms of AUC  with and without shortcut adjustment in the evaluation is shown in Figure 3a. For BASE-T  we observe that the shortcut adjustment w.r.t. alleles and peptide sources leads to an absolute AUC  decrease of 0.06 when applied separately, and a further decrease by 0.07 when combined, which indicates that the model learns shortcuts based on both MHC allele-specific and peptide source-specific features. In contrast, for ADA-T  the performance without adjustments and with source shortcut adjustment is nearly identical, which demonstrates that adversarial domain adaptation is effective in reducing shortcut learning based on peptide sources. When adjusting for both possible shortcuts in the evaluation, ADA-T leads to slightly reduced predictive performance compared to BASE-T, though.
@@
To verify that shortcuts are actually based on the label imbalance within different domains (see Figures 1a and 1b), we plot the prediction scores of peptides from the two most frequent peptide sources, Human betaherpesvirus 6B (mostly positives) and Vaccinia virus (mostly negatives) as histograms, separated according to the true response labels in Figure 3b. For BASE-T, we find that the distribution of prediction scores is always shifted towards the majority class within the peptide source: Human betaherpesvirus 6B peptides with a positive label tend to get higher prediction scores than Vaccinia virus peptides with a positive label. For negatives, the Vaccinia virus peptides get lower prediction scores than Human betaherpesvirus 6B peptides. ADA-T overcorrects the bias towards the majority label within sources for the T-cell response positives, as the Vaccinia virus peptides (mostly negatives) now get even higher prediction scores than Human betaherpesvirus 6B peptides. For the T-cell response negatives, the distributions of prediction scores corresponding to the two peptide sources are now well aligned, indicating that no source shortcuts are used. These findings may also explain the slight drop in overall prediction accuracy of ADA-T observed in Figure 3a, as the number of correct predictions in the minority classes are increased at the expense of additional errors in the majority classes.
@@
To test whether this is the case, we use t-stochastic neighborhood embedding (t-SNE) by Van der Maaten and Hinton (2008) to project the high-dimensional latent peptide representations $p$ to two-dimensional points. We here use the t-SNE implementation of the TensorFlow embedding projector with the default parameters for learning rate and perplexity.

### v2

@@ Abstract
We study the prediction of T-cell response for specific given peptides, which could, among other applications, be a crucial step towards the development of personalized cancer vaccines. It is a challenging task due to limited, heterogeneous training data featuring a multi-domain structure; such data entail the danger of shortcut learning, where models learn general characteristics of peptide sources, such as the source organism, rather than specific peptide characteristics associated with T-cell response. Using a transformer model for T-cell response prediction, we show that the danger of inflated predictive performance is not merely theoretical but occurs in practice. Consequently, we propose a domain-aware evaluation scheme. We then study different transfer learning techniques to deal with the multi-domain structure and shortcut learning. We demonstrate a per-source fine tuning approach to be effective across a wide range of peptide sources and further show that our final model is competitive with existing state-of-the-art approaches for predicting T-cell responses for human peptides.
keywords:
T cells, MHC, peptides, shortcut learning, domain adaptation, transformers
@@ 1 Introduction
The human immune system consists of humoral and cell-mediated mechanisms, with T cells playing a significant role in the latter. T cells recognize and eliminate cancerous or infected cells by detecting peptides presented on the cell surface by major histocompatibility complex (MHC) molecules, which can be grouped into two classes, MHC I and MHC II [1].
@@
Peptide-based vaccines are a promising approach for the personalized treatment of cancer [2] and might allow more precise vaccines against infectious diseases [3]. Such vaccines can be developed by selecting peptides that are specific to a pathogen or tumor. Since only a limited number of peptides can be included in the vaccine, their selection should be based on their probability of inducing a T-cell response. Predicting this probability is an important task in the development of peptide vaccines [2].
@@
From a computational perspective, this task can be separated into two sub-tasks: first predicting whether a peptide is presented on the cell surface by an MHC molecule and afterwards predicting whether a presented peptide induces a T-cell response. For the first task, a large amount of available experimental data has led to accurate machine learning-based predictions [4, 5, 6]. The second task can, despite some attempts [7, 8], still be considered as an open problem, in part due to fewer experimental data from T-cell assays [9].
@@
Aside from small sample sizes, another challenge arises from peptides in T-cell response data originating from different sources, such as viruses, bacteria, or human proteins. Moreover, peptides feature patterns that are specific to the allele of the MHC molecule they are presented by [10]. Both factors lead to data heterogeneity, which is either unaccounted for by existing approaches [8, 11] or handled by limiting the model to selected MHC alleles and considering small peptides, such that positions that are affected by MHC allele-specific patterns can be identified [12, 7, 13].
@@
To leverage all the available data, we investigate more flexible approaches that do not limit the selection of peptides based on their source, length, and presenting MHC allele. Based on our data analysis, we show how the heterogeneity of the resulting data set can be viewed as a multi-domain structure, which motivates the application of different transfer learning techniques in combination with the flexible transformer architecture [14]. We further investigate how the prediction performance on individual domains is affected by including data from the other domains in the training process, where improved performance is denoted as positive transfer and reduced performance as negative transfer [15].
@@
For all empirical studies it is critical to keep in mind that having several underlying domains with varying fractions of positives entails the risk of shortcut learning [16, 17, 18], that is, using domain-specific features to classify peptides based on their domain identity instead of T-cell response-specific patterns. To obtain shortcut-invariant model performance estimates, we thus propose a domain-aware evaluation scheme.
@@
Our results indicate that shortcuts based on peptide sources and MHC alleles are indeed learned by the transformer and lead to inflated performance estimates when the domain-aware evaluation is not applied. We further observe that per-source fine-tuning is effective in enabling positive transfer across a wide range of peptide sources and show that our final model is competitive with existing state-of-the-art approaches for predicting T-cell responses for human peptides.
@@ 2.1 Construction of a T-cell response data set
For some peptides in the data set, the MHC alleles are not specified in the IEDB. Additionally, the MHC allele information is incomplete since one peptide can often be presented on several MHC alleles. However, not all combinations of peptides and MHC alleles are tested. Since the existing MHC information in the IEDB is mostly based on predictions, we use existing models to obtain a list of MHC alleles for each peptide in a consistent way. NetMHCpan 4.1 [4] serves as a predictor for peptides with MHC class I alleles and NetMHCIIpan 4.0 [4] for peptides with MHC class II alleles. We select the 100 most frequent alleles, which account for 98.7% of peptide-MHC combinations in the IEDB data set, and use all MHC class-matched combinations of peptides and MHC alleles as input for the two predictors.
@@ 2.2 Domain structure in T-cell response data
Motivated by the size of the binding core in peptides [19], we define that peptides correspond to the same cluster if they share a sub-sequence of length nine. The Sankey diagram in Figure 1c confirms that the seven largest peptide clusters typically consist of peptides from only a few sources.
@@
The distinct binding motifs of MHC alleles [10] introduce an MHC-based domain structure in T-cell response data. We visualize the sequence logos for two exemplary MHC alleles and peptides of length nine, separately for T-cell response positives and negatives, in Figure 1d. At the anchor positions two and nine, there are obvious differences between the MHC alleles in the sequence logos. However, comparing sequence logos between T-cell response positives and negatives does not reveal obvious differences.
@@ 2.3 Transformer model
We use the transformer architecture [14], as shown in Figure 2, to capture T-cell response-specific patterns in peptide sequences. To apply the transformer, we formalize the data set to consist of data points $\{(\bm{x}_{i},y_{i})\}_{i=1}^{N_{\text{data}}}$, where $\bm{x}_{i}=(x_{i,1},\dots,x_{i,L})$ is the $i$-th peptide sequence with each of the $x_{i,l}$ representing one of 20 amino acids or a padding token. The padding token allows the sequences to all have the same length $L$, which corresponds to the maximal peptide length. We denote the binary label indicating whether the peptide $\bm{x}_{i}$ leads to a T-cell response with $y_{i}$.
Our model gets a one-hot representation of the peptide sequence $\bm{x}_{i}$ as input. Then the model maps each amino acid or padding token $x_{i,j}$ of the peptide sequence to a learned embedding $\bm{e}_{i,l}\in\mathbb{R}^{d/2}$. To tell the transformer the positions at which the amino acids occur in the peptide, we define a learned embedding $\bm{p}_{l}\in\mathbb{R}^{d/2}$ for each position $l\in\{1,\dots,L\}$. Additionally, we use positional encodings $\bm{s}_{l}\in\mathbb{R}^{d/2}$ based on sine functions with varying frequencies (see [14] for more details). We then add the two variants of positional vectors and concatenate them with the amino acid embeddings to obtain a vector representation $\left(\begin{smallmatrix}\bm{e}_{i,l}\\ \bm{p}_{l}+\bm{s}_{l}\end{smallmatrix}\right)\in\mathbb{R}^{d}$ for each element $x_{i,l}$ of the peptide sequence $\bm{x}_{i}$.
The resulting sequence of vector representations is the input of a transformer encoder layer, which updates the vector representations based on interactions between amino acids using a mechanism called multi-head self-attention (see [14] for more details).
As the output of the transformer encoder is again a sequence of vector representations, we apply mean pooling to obtain a single vector $\bm{h}_{i}\in\mathbb{R}^{d}$, which is a latent representation of the $i$-th peptide. In a similar way as described here, the transformer architecture has been used in earlier work to learn representations of peptides [6] and proteins [20].
@@ 2.4 Adversarial domain adaptation
To address these issues, we apply a transfer learning technique called adversarial domain adaptation [21], to encourage the model to perform predictions independently of the domain identity of a peptide. Unlike the typical domain adaptation setting with a large amount of training data for a source domain and no or little training data for the target domain [22], in our setting, there are more than two domains and all domains can act as target and source domain at the same time. We use a variant of adversarial domain adaptation that is adapted to our setting. This variant is very similar to multi-domain adversarial learning [23].
@@ 2.5 Per-source fine-tuning
When one machine learning model is trained on data from several domains, the performance on individual domains can be reduced compared to models that are trained on data from one domain only. Such instances of negative transfer are more likely to occur if domains are too dissimilar [15]. Negative transfer is also possible when applying adversarial domain adaptation because there is only one model for all domains. Additionally, data points from all domains are encouraged to be represented similarly in the model. To avoid negative transfer, a transfer learning technique is needed that provides more flexibility to adapt to individual domains, while still allowing positive transfer from other domains.
As an alternative to adversarial domain adaptation, we propose to pre-train a T-cell response model until convergence on all domains and then continue training on only one domain. This second step is also called fine-tuning. We allow all model parameters to be updated during fine-tuning such that the model has the necessary flexibility to correct instances of negative transfer. This transfer learning technique is commonly applied in combination with the transformer and proteins [20] or peptides [6].
@@ 2.6 Bag of amino acids baseline
To allow a direct comparison of our proposed transformer-based models to a simpler model that is trained on the same data set, we use a bag of amino acids (Bag-Of-AA) baseline, since it can be viewed as a simplified version of PRIME [7].
This baseline model counts how often each of the 20 amino acids occurs in a given peptide and puts the frequencies into a 20-dimensional vector. An MLP with one hidden layer receives these vectors as input and is trained to predict the T-cell responses. As with the transformer models, we use the binary cross-entropy as training objective. The Bag-Of-AA model does not consider positional information, that is, it does not see at which positions the amino acids occur in the peptide. The amino acid frequencies allow the model to capture possible preferences for certain amino acids that T cells might have.
@@ 2.7 Model evaluation
Within a peptide cluster, sequences can be nearly identical (cf. Section 2.2). Splitting the data completely at random could lead to such near identical sequences being distributed among training, validation, and test sets. This can potentially cause inflated generalization estimates. To prevent this, we assign peptide clusters as a whole (instead of single peptides) to a randomly selected partition [19].
@@ 2.8 Detection of shortcut learning
In our analysis, we evaluate model performance to account for such shortcuts. In particular, we group the evaluation data into subsets such that each subset contains peptides from only one combination of peptide source and MHC allele. We further weight by subset size and peptide:MHC assignments to ensure each peptide contributes equally to the overall performance (see Supplement S2 for details). Since this evaluation setting ensures that performance estimates are not inflated by the use of shortcuts, we refer to this strategy as shortcut adjusted evaluation.
@@ 3 Results
Figure 3: Shortcut learning and the effect of adversarial domain adaptation. The left column shows results for BASE-T and the right column shows corresponding results for ADA-T with adversarial domain adaptation being applied on peptide sources. (a) Model performance on validation data with different settings of accounting for shortcuts in the evaluation. For the ”allele adjusted” performance, peptides are grouped by MHC alleles in the evaluation to detect shortcuts based on MHC alleles. ”Source adjusted” is analogous for grouping peptides by their source. (b) Distribution of prediction scores, separated by the T-cell response labels. For each label, two distributions are shown that correspond to the two most frequent peptide sources, Human betaherpesvirus 6B (majority of labels is positive) and Vaccinia virus (majority of labels is negative). The markers on the x-axis show the mean prediction scores of the two distributions. (c) t-SNE visualization of the latent peptide representations $\mathbf{h}$.
@@ 3.1 Shortcut learning in practice
We investigate the prevalence of shortcut learning in T-cell response prediction models trained on the full T-cell response data set, that is, on all peptide sources and MHC alleles (MHC I+II) combined. In order to study whether shortcut learning can be prevented by encouraging domain invariance in the model architecture and whether this constraint could serve as regularizer to improve the model performance, we compare two different models: the baseline transformer BASE-T, and ADA-T, the adversarial domain adaptation approach with domain adaptation strength set to $\lambda=10$. The results for both models are summarized side by side in Figure 3.
@@
The prediction performance in terms of AUC  with and without shortcut adjustment in the evaluation is shown in Figure 3a. For BASE-T  we observe that the shortcut adjustment w.r.t. alleles and peptide sources leads to an absolute AUC  decrease of 0.06 when applied separately, and a further decrease by 0.07 when combined, which indicates that the model learns shortcuts based on both MHC allele-specific and peptide source-specific features. In contrast, for ADA-T  the performance without adjustments and with source shortcut adjustment is nearly identical, which demonstrates that adversarial domain adaptation is effective in reducing shortcut learning based on peptide sources. When adjusting for both possible shortcuts in the evaluation, ADA-T leads to slightly reduced predictive performance compared to BASE-T.
@@
To verify that shortcuts are actually based on the label imbalance within different domains (see Figures 1a and 1b), we plot the prediction scores of peptides from the two most frequent peptide sources, Human betaherpesvirus 6B (mostly positives) and Vaccinia virus (mostly negatives) as histograms, separated according to the true response labels in Figure 3b. For BASE-T, we find that the distribution of prediction scores is always shifted towards the majority class within the peptide source: Human betaherpesvirus 6B peptides with a positive label tend to get higher prediction scores than Vaccinia virus peptides with a positive label. For negatives, the Vaccinia virus peptides get lower prediction scores than Human betaherpesvirus 6B peptides. ADA-T overcorrects the bias towards the majority label within sources for the T-cell response positives, as the Vaccinia virus peptides (mostly negatives) now get even higher prediction scores than Human betaherpesvirus 6B peptides. For the T-cell response negatives, the distributions of prediction scores corresponding to the two peptide sources are now well aligned, indicating that no source shortcuts are used. These findings may also explain the slight drop in overall prediction accuracy of ADA-T observed in Figure 3a, as the number of correct predictions in the minority classes are increased at the expense of additional errors in the majority classes. We also carried out this study for other peptide sources (Supplement S4), where a similar effect is visible for most of them, albeit to a smaller degree, which is likely due to the smaller sample size.
@@
To test whether this is the case, we use t-stochastic neighborhood embedding (t-SNE) by [24] to project the high-dimensional latent peptide representations $p$ to two-dimensional points. We here use the t-SNE implementation of the TensorFlow embedding projector with the default parameters for learning rate and perplexity.